# MM-UNet: Morph Mamba U-shaped Convolutional Networks for Retinal Vessel Segmentation

### v1

@@ I Introduction
In recent years, the growing demand for early diagnosis of retinal fundus diseases has drawn increasing attention to automated retinal vessel analysis [1]. With advances in imaging techniques and computational hardware, Deep Learning (DL) based methods [18, 25, 20] have emerged as the dominant approach for Retinal Vessel Segmentation (RVS). These methods offer high precision in vascular abnormalities, providing critical support for disease screening, diagnostic efficiency, and streamlined clinical workflows [35].
@@
Unlike lesion [16, 14, 40] or organ segmentation [33, 24], retinal-vessel segmentation confronts a precision bottleneck [27, 17]. The main culprit is the vessels’ highly intricate tubular topology and complex clinical imaging. As shown in Fig. 1 (a), retinal vessels possess extremely thin terminal branches and display significant global morphological deformation. This results in vessel targets occupying only a small portion of the local pixel space in retinal images, which poses a serious challenge for deep learning models built on convolutional operations. Those models, such as U-Net [30] and its variants [2, 15], rely heavily on standard convolutional upsampling operations, limiting their ability to precisely capture high-resolution peripheral details and frequently leading to fragmented or disconnected segmentation outputs. With the emergence of Vision Transformer models in the field of retinal vessel segmentation [34, 9, 24], global feature extraction capabilities have been partially improved. However, these methods still lack the capacity to provide precise guidance for capturing delicate local patterns, particularly at vessel termini [27]. Recent methods such as DCASU-Net [37], FR-UNet [20], and FSG-Net-L [31] have improved local representation using multi-scale fusion and high-resolution strategies, yet accurately capturing the full vascular topology remains challenging.
@@
To tackle the above obstacles, this work proposes MM-UNet (Morph Mamba U-shaped convolutional Networks), a novel segmentation framework featuring two innovative components: MMC (Morph Mamba Convolution) layers and RSSG (Reverse Selective State Guidance) module. Specifically, as shown in Fig. 1 (b), the MMC layers tackle the intricate topological tubular structure of retinal vessels by integrating a dynamic morph convolution mechanism with morph state-space modeling. This integration enables the adaptive capture of narrow and tortuous local features inherent in tubular structures, significantly enhancing geometric perception. By superseding pointwise convolutional layers within all of the U-shaped segmentation networks, MMC significantly enhances the model’s capability for efficient topological representation and modeling. Concurrently, RSSG modules are introduced into the inter-level skip connections of the U-shaped architecture to provide additional geometric structural guidance during the upsampling stage. Through the combination of reverse guidance mechanisms and state-space modeling, RSSG modules extract complementary boundary information from both interior and exterior regions encoded during downsampling, effectively enhancing the model’s capacity to discern structural contours and maintain spatial coherence. Extensive experiments conducted on two widely used retinal image datasets, STARE and DRIVE, demonstrate the superior performance of MM-UNet and its key components compared to current state-of-the-art methods, achieving improvements of at least 1.64% and 1.25% in terms of F1-score, respectively. In summary, our contributions are as follows:
@@ II Relate Works
Convolutional Segmentation Networks mark the earliest deep learning solutions in this domain. U-Net [30] pioneered the encoder–decoder paradigm with skip connections, enabling effective recovery of spatial details. However, standard convolutional methods often struggle with preserving the global continuity of thin, tortuous vessels, especially in regions with sparse signals.
Graph-Based and Multi-Scale Hybrid Models aim to better capture vessel topology and multi-scale feature representation. DE-DCGCN-EE [11] employs dynamic graph convolution and edge enhancement to model vessel connectivity, while GT-DLA-dsHFF [13] integrates global and local attention via deep–shallow hierarchical fusion. BCU-Net [39] and PA-Net [22] further combine multi-resolution encoding with adaptive fusion to improve detail recognition. Despite their structural modeling capabilities, these methods often exhibit limited boundary localization, leading to suboptimal F1 and sensitivity scores.
Transformer and Attention-Enhanced Architectures improve long-range dependency modeling, which is essential for maintaining vessel continuity. For example, Wave‑Net [21] replaces U‑Net’s standard skip‑connections with a detail‑enhancement-and-denoising block to better preserve continuity even in ultra-thin vessel branches. CoVi‑Net [8] integrates a local‑global feature aggregation module and employs bidirectional weighted fusion to reinforce structural consistency across the segmented vasculature. However, this series of Transformer-based methods tends to suffer from overfitting during training and exhibits limited sensitivity to fine vascular branches [27].
State-Space or Dual-Decoder Frameworks have recently emerged to balance modeling capacity with computational efficiency. MSPDD-Net [10] adopts dual decoders and wavelet edge enhancement, but its static context still limits local structural flexibility. HRD-Net [19] utilizes deformable convolutions in high-resolution pipelines to preserve microvascular details, yet may lack adaptability in sparse or irregular vessel regions. TP-Net [28] introduces a two-path design to decouple edge and trunk extraction, but may suffer from coarse-to-fine fusion inconsistencies.
@@ III-A U-shaped Framework Overview
MM-UNet adopts a U-shaped segmentation framework, which is shown in Fig. 2 (a), comprising a five-step encoder and a corresponding four-step decoder. All pointwise convolutional layers are replaced by the proposed Morph Mamba Convolution (MMC) layers (detailed in Section III-B). The encoder is adapted from non-pretraining ResNet-34 [6] to extract five downsampled features $F_{i}$ with dimensions $\frac{H}{2^{i}}\times\frac{W}{2^{i}}\times C_{i}$, where $C_{1}=64$ and $C_{i}=64\times 2^{i-2}$, $i\in[2,3,4,5]$, from input images $x$ of size $H\times W\times 3$. To reduce computational cost [3], three MMC layers are used to unify the channels of $F_{3}$, $F_{4}$, and $F_{5}$ to 64 (defined as $F_{3}^{\prime}$, $F_{4}^{\prime}$, and $F_{5}^{\prime}$). After background suppression by the Convolutional Block Attention Module (CBAM) [36], the low-level feature $F_{1}$ is passed to the proposed Reverse Selective State Guidance (RSSG) modules (detailed in Section III-C) and integrated into the decoder to guide upsampling and enhance geometric boundary awareness. Finally, multi-scale decoder features are fused and passed through a sigmoid layer to generate a pixel-wise probability map for accurate and scalable retinal vessel segmentation.
@@ III-B Morph Mamba Convolution Layers
To overcome the limited sensitivity of traditional pointwise convolutions to complex branching topologies [38], the MMC layer integrates morph convolution with a morph state-space computing mechanism, as illustrated in Fig. 2 (b). Assume a conventional $3\times 3$ kernel-based convolution layer with the convolution center located at $C_{i}=\{x_{i},y_{i}\}$. Thus, the other coordinates within its receptive field can be expressed as in Eq. 1a. In MMC, a learnable morph offset $\Delta\in\{-1,1\}$ is introduced to redefine the surrounding positions relative to the center, yielding $C_{i+c}=\{x_{i+c},y_{i+c}\}$ and $C_{j+c}=\{x_{j+c},y_{j+c}\}$, as shown in Eq. 1b,  1c. Due to the fractional nature of $\Delta$, sub-pixel level displacements are allowed, and the corresponding feature values at non-integer positions are obtained via bilinear interpolation, as shown in Eq. 1d.
@@
To enhance the model’s perception of complex tubular structures, MMC incorporates a Multi-view Feature Fusion Strategy [26] along with a selective state-space computation mechanism [5], guiding the model to complement its attention to fundamental features from multiple perspectives. Specifically, for each position $C$, two feature maps $f^{l}(C_{x})$ and $f^{l}(C_{y})$ are extracted along the x-axis and y-axis, respectively, as computed in Eq. 2a. Meanwhile, based on a novel morph SSM computation, cross-axis contextual dependencies are constructed, as illustrated in Eq. 2b. Finally, as shown in Eq. 2c, the feature sampling operation is completed by stacking multi-view feature maps from different perspectives to form a correlated multi-view representation.
@@ III-C Reverse Selective State Guidance
Motivated by the blurred boundary representations often observed in deep-sampled features [23], the RSSG module is designed to provide effective geometric structure guidance during upsampling. As illustrated in Fig. 2 (c), it integrates reverse guidance theory with the state-space mechanism to enhance the model’s ability to perceive geometric boundaries, thereby improving its focus on complex vascular topologies.
Specifically, each RSSG module takes three inputs: the shallow edge feature $F_{edge}$ extracted by CBAM, the deep semantic feature $F_{i}^{\prime}$ obtained from the corresponding downsampling stage, and the reverse prediction $R_{i}$ from the preceding decoder block. As shown in Eq. 3, guided by the detailed spatial information from shallow features [4], the RSSG module constructs a unified feature state-space by integrating decoder features and semantic base representations, which enhances the decoding process and promotes more accurate segmentation around geometrically complex boundaries.
@@
where $out_{r}$ denotes the output of the RSSG module, mlpmlp represents a multilayer perceptron, $mmc(\cdot)$ indicates an MMC layer with a $3\times 3$ kernel, and Sigmoid denotes the sigmoid activation function [29].
@@ IV-A Datasets
DRIVE [32]: This dataset consists of 40 fundus images (565 × 584 pixels) collected from a diabetic retinopathy screening program in the Netherlands. The images are evenly split into training and test sets. Each image includes a field-of-view mask and corresponding vessel annotations. For the test set, two sets of vessel labels are provided: one as the gold standard and one from a second human observer. To facilitate model training, all images are resized to 608 × 608 pixels.
STARE [7]: This dataset contains 20 fundus images, each with a resolution of 605 × 700 pixels. Half of the cases present retinal vascular abnormalities. All images are annotated by clinical experts. Fifty percent of the images are allocated for training, and the remainder for testing. To facilitate model training, all images are resized to 704 × 704 pixels.
@@ IV-B Implementation Details
   DRIVE    U-Net [30]    95.56    75.56    97.30    79.97
   DE-DCGCN-EE [11]    97.05    83.59    98.26    82.88
@@
   BCU-Net [39]    96.62    82.38    98.00    80.89
   Wave-Net [21]    95.61    81.64    97.64    82.54
   CoVi-Net [8]    96.98    83.47    98.30    87.48
   HRD-Net [19]    97.04    83.71    98.33    83.12
   PA-Net [22]    95.82    82.84    98.07    83.93
   MSPDD-Net [10]    97.45    87.51    98.21    87.95
@@
   STARE    U-Net [30]    96.17    81.67    98.33    81.12
   DE-DCGCN-EE [11]    97.51    84.05    98.61    83.63
@@
   BCU-Net [39]    97.01    85.00    98.07    82.23
   Wave-Net [21]    96.41    79.02    98.36    81.40
   CoVi-Net [8]    97.61    83.05    98.87    90.31
   HRD-Net [19]    97.55    84.59    98.62    83.57
   PA-Net [22]    97.09    88.13    98.05    85.61
   MSPDD-Net [10]    97.76    88.75    98.91    90.52
@@ IV-C Performance
  • Traditional Convolutional Segmentation Methods: U-shaped Convolutional Network (U-Net) [30].
  • Graph and Multi-scale Convolutional Models: Dual Encoder-based Dynamic-channel Graph Convolutional Network with Edge Enhancement (DE-DCGCN-EE) [11], Global Transformer and Dual Local Attention Network via Deep-Shallow Hierarchical Feature Fusion (GT-DLA-dsHFF) [13], Bridge ConvNeXt U-Net (BCU-Net) [39], and a Hybrid Architecture based on LPT and AFFM (PA-Net)  [22].
  • Transformer and Attention-enhanced Models: Wave-Net [21] and Convolutional Vision Transformer Network (CoVi-Net) [8].
  • State-space and Dual-decoder Architectures: Mamba Semantic Perception Dual-decoding Network (MSPDD-Net) [10], High Resolution based on Deformable Convolution v3 (HRD-Net) [19], and Two-Path Network (TP-Net) [28].

### v2

@@ I Introduction
In recent years, the growing demand for early diagnosis of retinal fundus diseases has drawn increasing attention to automated retinal vessel analysis [cen2021automatic]. With advances in imaging techniques and computational hardware, Deep Learning (DL) based methods [liang2025sfit, pang2025slim, liu2022full] have emerged as the dominant approach for Retinal Vessel Segmentation (RVS). These methods offer high precision in vascular abnormalities, providing critical support for disease screening, diagnostic efficiency, and streamlined clinical workflows [wang2024advancing].
@@
Unlike lesion [liang2024comprehensive, liang2025hwa, zhou2023ga] or organ segmentation [tadokoro2023pre, pang2023slim], retinal-vessel segmentation confronts a precision bottleneck [qin2024review, liang2022u]. The main culprit is the vessels’ highly intricate tubular topology and complex clinical imaging. As shown in Fig. 1 (a), retinal vessels possess extremely thin terminal branches and display significant global morphological deformation. This results in vessel targets occupying only a small portion of the local pixel space in retinal images, which poses a serious challenge for deep learning models built on convolutional operations. Those models, such as U-Net [ronneberger2015u] and its variants [dai2024i2u, liang2023agilenet], rely heavily on standard convolutional upsampling operations, limiting their ability to precisely capture high-resolution peripheral details and frequently leading to fragmented or disconnected segmentation outputs. With the emergence of Vision Transformer models in the field of retinal vessel segmentation [tan2024deep, kreitner2024synthetic, pang2023slim], global feature extraction capabilities have been partially improved. However, these methods still lack the capacity to provide precise guidance for capturing delicate local patterns, particularly at vessel termini [qin2024review]. Recent methods such as DCASU-Net [xu2023dcsau], FR-UNet [liu2022full], and FSG-Net-L [seo2025full] have improved local representation using multi-scale fusion and high-resolution strategies, yet accurately capturing the full vascular topology remains challenging.
@@
To tackle the above obstacles, this work proposes MM-UNet (Morph Mamba U-shaped convolutional Networks), a novel segmentation framework featuring two innovative components: MMC (Morph Mamba Convolution) layers and RSSG (Reverse Selective State Guidance) module. Specifically, as shown in Fig. 1 (b), the MMC layers tackle the intricate topological tubular structure of retinal vessels by integrating a dynamic morph convolution mechanism with morph state-space modeling. This integration enables the adaptive capture of narrow and tortuous local features inherent in tubular structures, significantly enhancing geometric perception. By superseding pointwise convolutional layers within all of the U-shaped segmentation networks, MMC significantly enhances the model’s capability for efficient topological representation and modeling. Concurrently, RSSG modules are introduced into the inter-level skip connections of the U-shaped architecture to provide additional geometric structural guidance during the upsampling stage. Through the combination of reverse guidance mechanisms and state-space modeling, RSSG modules extract complementary boundary information from both interior and exterior regions encoded during downsampling, effectively enhancing the model’s capacity to discern structural contours and maintain spatial coherence. Extensive experiments conducted on two widely used retinal image datasets, STARE and DRIVE, demonstrate the superior performance of MM-UNet and its key components compared to current state-of-the-art methods, achieving improvements of at least 1.64% and 1.25% in terms of F1-score, respectively. In summary, our contributions are as follows:These authors contributed equally to this work
@@ II Relate Works
Convolutional Segmentation Networks mark the earliest deep learning solutions in this domain. U-Net [ronneberger2015u] pioneered the encoder–decoder paradigm with skip connections, enabling effective recovery of spatial details. However, standard convolutional methods often struggle with preserving the global continuity of thin, tortuous vessels, especially in regions with sparse signals.
Graph-Based and Multi-Scale Hybrid Models aim to better capture vessel topology and multi-scale feature representation. DE-DCGCN-EE [li2022dual] employs dynamic graph convolution and edge enhancement to model vessel connectivity, while GT-DLA-dsHFF [s22155826] integrates global and local attention via deep–shallow hierarchical fusion. BCU-Net [ZHANG2023106960] and PA-Net [luo2025pa] further combine multi-resolution encoding with adaptive fusion to improve detail recognition. Despite their structural modeling capabilities, these methods often exhibit limited boundary localization, leading to suboptimal F1 and sensitivity scores.
Transformer and Attention-Enhanced Architectures improve long-range dependency modeling, which is essential for maintaining vessel continuity. For example, Wave‑Net [LIU2023106341] replaces U‑Net’s standard skip‑connections with a detail‑enhancement-and-denoising block to better preserve continuity even in ultra-thin vessel branches. CoVi‑Net [JIANG2024108047] integrates a local‑global feature aggregation module and employs bidirectional weighted fusion to reinforce structural consistency across the segmented vasculature. However, this series of Transformer-based methods tends to suffer from overfitting during training and exhibits limited sensitivity to fine vascular branches [qin2024review].
State-Space or Dual-Decoder Frameworks have recently emerged to balance modeling capacity with computational efficiency. MSPDD-Net [li2025mspdd] adopts dual decoders and wavelet edge enhancement, but its static context still limits local structural flexibility. HRD-Net [LIU2024108295] utilizes deformable convolutions in high-resolution pipelines to preserve microvascular details, yet may lack adaptability in sparse or irregular vessel regions. TP-Net [10018415] introduces a two-path design to decouple edge and trunk extraction, but may suffer from coarse-to-fine fusion inconsistencies.
@@ III-A U-shaped Framework Overview
MM-UNet adopts a U-shaped segmentation framework, which is shown in Fig. 2 (a), comprising a five-step encoder and a corresponding four-step decoder. All pointwise convolutional layers are replaced by the proposed Morph Mamba Convolution (MMC) layers (detailed in Section III-B). The encoder is adapted from non-pretraining ResNet-34 [he2016deep] to extract five downsampled features $F_{i}$ with dimensions $\frac{H}{2^{i}}\times\frac{W}{2^{i}}\times C_{i}$, where $C_{1}=64$ and $C_{i}=64\times 2^{i-2}$, $i\in[2,3,4,5]$, from input images $x$ of size $H\times W\times 3$. To reduce computational cost [du2025net], three MMC layers are used to unify the channels of $F_{3}$, $F_{4}$, and $F_{5}$ to 64 (defined as $F_{3}^{\prime}$, $F_{4}^{\prime}$, and $F_{5}^{\prime}$). After background suppression by the Convolutional Block Attention Module (CBAM) [woo2018cbam], the low-level feature $F_{1}$ is passed to the proposed Reverse Selective State Guidance (RSSG) modules (detailed in Section III-C) and integrated into the decoder to guide upsampling and enhance geometric boundary awareness. Finally, multi-scale decoder features are fused and passed through a sigmoid layer to generate a pixel-wise probability map for accurate and scalable retinal vessel segmentation.
@@ III-B Morph Mamba Convolution Layers
To overcome the limited sensitivity of traditional pointwise convolutions to complex branching topologies [yang2024d], the MMC layer integrates morph convolution with a morph state-space computing mechanism, as illustrated in Fig. 2 (b). Assume a conventional $3\times 3$ kernel-based convolution layer with the convolution center located at $C_{i}=\{x_{i},y_{i}\}$. Thus, the other coordinates within its receptive field can be expressed as in Eq. 1a. In MMC, a learnable morph offset $\Delta\in\{-1,1\}$ is introduced to redefine the surrounding positions relative to the center, yielding $C_{i+c}=\{x_{i+c},y_{i+c}\}$ and $C_{j+c}=\{x_{j+c},y_{j+c}\}$, as shown in Eq. 1b,  1c. Due to the fractional nature of $\Delta$, sub-pixel level displacements are allowed, and the corresponding feature values at non-integer positions are obtained via bilinear interpolation, as shown in Eq. 1d.
@@
To enhance the model’s perception of complex tubular structures, MMC incorporates a Multi-view Feature Fusion Strategy [qi2023dynamic] along with a selective state-space computation mechanism [gu2023mamba], guiding the model to complement its attention to fundamental features from multiple perspectives. Specifically, for each position $C$, two feature maps $f^{l}(C_{x})$ and $f^{l}(C_{y})$ are extracted along the x-axis and y-axis, respectively, as computed in Eq. 2a. Meanwhile, based on a novel morph SSM computation, cross-axis contextual dependencies are constructed, as illustrated in Eq. 2b. Finally, as shown in Eq. 2c, the feature sampling operation is completed by stacking multi-view feature maps from different perspectives to form a correlated multi-view representation.
@@ III-C Reverse Selective State Guidance
Motivated by the blurred boundary representations often observed in deep-sampled features [pang2025efficient], the RSSG module is designed to provide effective geometric structure guidance during upsampling. As illustrated in Fig. 2 (c), it integrates reverse guidance theory with the state-space mechanism to enhance the model’s ability to perceive geometric boundaries, thereby improving its focus on complex vascular topologies.
Specifically, each RSSG module takes three inputs: the shallow edge feature $F_{edge}$ extracted by CBAM, the deep semantic feature $F_{i}^{\prime}$ obtained from the corresponding downsampling stage, and the reverse prediction $R_{i}$ from the preceding decoder block. As shown in Eq. 3, guided by the detailed spatial information from shallow features [du2022icgnet], the RSSG module constructs a unified feature state-space by integrating decoder features and semantic base representations, which enhances the decoding process and promotes more accurate segmentation around geometrically complex boundaries.
@@
where $out_{r}$ denotes the output of the RSSG module, mlpmlp represents a multilayer perceptron, $mmc(\cdot)$ indicates an MMC layer with a $3\times 3$ kernel, and Sigmoid denotes the sigmoid activation function [ren2007potential].
@@ IV-A Datasets
DRIVE [staal2004ridge]: This dataset consists of 40 fundus images (565 × 584 pixels) collected from a diabetic retinopathy screening program in the Netherlands. The images are evenly split into training and test sets. Each image includes a field-of-view mask and corresponding vessel annotations. For the test set, two sets of vessel labels are provided: one as the gold standard and one from a second human observer. To facilitate model training, all images are resized to 608 × 608 pixels.
STARE [hoover2000locating]: This dataset contains 20 fundus images, each with a resolution of 605 × 700 pixels. Half of the cases present retinal vascular abnormalities. All images are annotated by clinical experts. Fifty percent of the images are allocated for training, and the remainder for testing. To facilitate model training, all images are resized to 704 × 704 pixels.
@@ IV-B Implementation Details
   DRIVE    U-Net [ronneberger2015u]    95.56    75.56    97.30    79.97
   DE-DCGCN-EE [li2022dual]    97.05    83.59    98.26    82.88
@@
   BCU-Net [ZHANG2023106960]    96.62    82.38    98.00    80.89
   Wave-Net [LIU2023106341]    95.61    81.64    97.64    82.54
   CoVi-Net [JIANG2024108047]    96.98    83.47    98.30    87.48
   HRD-Net [LIU2024108295]    97.04    83.71    98.33    83.12
   PA-Net [luo2025pa]    95.82    82.84    98.07    83.93
   MSPDD-Net [li2025mspdd]    97.45    87.51    98.21    87.95
@@
   STARE    U-Net [ronneberger2015u]    96.17    81.67    98.33    81.12
   DE-DCGCN-EE [li2022dual]    97.51    84.05    98.61    83.63
@@
   BCU-Net [ZHANG2023106960]    97.01    85.00    98.07    82.23
   Wave-Net [LIU2023106341]    96.41    79.02    98.36    81.40
   CoVi-Net [JIANG2024108047]    97.61    83.05    98.87    90.31
   HRD-Net [LIU2024108295]    97.55    84.59    98.62    83.57
   PA-Net [luo2025pa]    97.09    88.13    98.05    85.61
   MSPDD-Net [li2025mspdd]    97.76    88.75    98.91    90.52
@@ IV-C Performance
  • Traditional Convolutional Segmentation Methods: U-shaped Convolutional Network (U-Net) [ronneberger2015u].
  • Graph and Multi-scale Convolutional Models: Dual Encoder-based Dynamic-channel Graph Convolutional Network with Edge Enhancement (DE-DCGCN-EE) [li2022dual], Global Transformer and Dual Local Attention Network via Deep-Shallow Hierarchical Feature Fusion (GT-DLA-dsHFF) [s22155826], Bridge ConvNeXt U-Net (BCU-Net) [ZHANG2023106960], and a Hybrid Architecture based on LPT and AFFM (PA-Net)  [luo2025pa].
  • Transformer and Attention-enhanced Models: Wave-Net [LIU2023106341] and Convolutional Vision Transformer Network (CoVi-Net) [JIANG2024108047].
  • State-space and Dual-decoder Architectures: Mamba Semantic Perception Dual-decoding Network (MSPDD-Net) [li2025mspdd], High Resolution based on Deformable Convolution v3 (HRD-Net) [LIU2024108295], and Two-Path Network (TP-Net) [10018415].